13. Write the equation of the line which has y-intercept (0, 5) and is perpendicular to the line with equation y = –3x + 1. Write the equation in slope-intercept form.
Answer by rodriguezh2(15)

You must use the equation (y - y1)= m(x -x1)to solve this problem.
We know that the line we are looking for is perpendicular to the line
y = -3x + 1.
Perpendicular lines have negative reciprocal slopes so our
slope(m) must equal 1/3.
We also know that point (0, 5) is the y intercept.
Now we plug into the equation:
(y - 5) = (1/3)(x - 0)
(y - 5) = (1/3)x + 0
y = (1/3)x + 5 is the equation of the line